# Non-Orthogonal Affine Frequency Division Multiplexing for Spectrally Efficient High-Mobility Communications

### v1

@@ Abstract
This paper proposes a novel non-orthogonal affine frequency division multiplexing (nAFDM) waveform for reliable high-mobility communications with enhanced spectral efficiency (SE). The key idea is to introduce a bandwidth compression factor into the AFDM modulator to enable controllable subcarrier overlapping. We first detail the proposed nAFDM transceiver and derive the corresponding input-output signal relationship. Then, an efficient nAFDM signal generation method based on the inverse discrete Fourier transform (IDFT) is proposed, enabling practical implementation using existing inverse fast Fourier transform (IFFT) modules without additional hardware complexity. Next, to characterize the impact of non-orthogonal modulation, we derive a closed-form expression of inter-carrier interference (ICI), showing its dependence on the bandwidth compression factor. To mitigate the resulting interference, we propose a soft iterative detection algorithm and a low-complexity implementation approach that leverages the distribution characteristics of ICI. Simulation results demonstrate that 1) in terms of bit error rate (BER), the proposed nAFDM can achieve near identical BER compared to conventional AFDM, while outperforms other waveform counterparts; 2) nAFDM is capable of striking higher SE compared to other existing waveforms; and 3) the proposed nAFDM achieves an attractive BER vs. SE trade-off, and the proposed soft ID scheme can attain a trade-off between BER and complexity.
Index Terms:
Affine frequency division multiplexing (AFDM), doubly-selective channels, inter-carrier interference (ICI), non-orthogonal waveform, spectrally efficient frequency division multiplexing (SEFDM), soft iterative detection.
@@ I-A Background
Providing reliable wireless connectivity in high-mobility scenarios, such as vehicle-to-everything systems, high-speed railways, unmanned aerial vehicles, and low-earth-orbit satellite networks, is crucial for the upcoming sixth generation systems [1, 2]. The widely adopted orthogonal frequency division multiplexing (OFDM) systems may suffer from severe inter-carrier interference (ICI) in high-mobility environments due to large Doppler shifts, leading to significant degradation in communication performance [3]. Against this problem, affine frequency division multiplexing (AFDM) has emerged recently as an excellent chirp-based multicarrier modulation scheme that offers improved reliability in high-mobility channels [4, 5]. In AFDM, data symbols are modulated over multiple orthogonal chirp subcarriers, thus spreading each information symbol across the entire bandwidth. Such a modulation is carried out in inverse discrete affine Fourier transform (IDAFT) [6], which can be efficiently implemented with the aid of inverse fast Fourier transform (IFFT) [5]. At the receiver side, the DAFT is applied to recover the transmitted symbols. Since DAFT can be regarded as a generalization of both the discrete Fourier transform (DFT) and the discrete Fresnel transform (DFnT) [7], AFDM can be built upon the minimum modification of OFDM systems and includes orthogonal chirp division multiplexing (OCDM) as a special case [8]. Compared to OFDM and OCDM, AFDM enables separable channel representation and achieves full diversity in doubly selective channels by appropriately tuning the chirp rate to match the Doppler profile [5, 9]. These advantages make AFDM an attractive waveform for enhanced communication performance in high-mobility environments.
@@ I-B Related Works
Several detection and channel estimation algorithms have been proposed to enhance the performance of AFDM. These include a maximum ratio combining scheme [5], an iterative minimum mean square error (MMSE) receiver that balances complexity and detection reliability [9], and a message passing algorithm that exploits the sparse structure of the DAFT-domain channel to jointly perform interference cancellation and signal detection with reduced complexity [10]. Pilot-aided channel estimation for AFDM systems was first introduced in [5], where pilot symbols are embedded into the transmitted frame to efficiently estimate channel parameters. This approach was later extended to multiple-input multiple-output (MIMO)-AFDM systems in [11], where a low-complexity diagonal reconstruction algorithm was developed to estimate the effective channel matrix. To enhance spectral efficiency (SE), a channel estimation scheme leveraging superimposed pilots in the DAFT domain has been proposed [12].
Besides, AFDM has been integrated with other advanced transmission techniques for a wide range of applications. For example, it has been introduced to serve as the foundational multicarrier waveform to support massive high-mobility machine-type communications through the integration of sparse code multiple access [13]. In addition, AFDM with index modulation [14, 15] and generalized spatial modulation aided AFDM [16] are proposed for improved bit error rate (BER) performance and energy efficiency. Furthermore, AFDM has been employed in integrated sensing and communications to enable accurate sensing and reliable parameter estimation [17, 18, 19].
@@
To meet the ever-growing demand for higher SE, non-orthogonal multicarrier waveform designs have attracted increasing research attention. One notable example is spectrally efficient frequency division multiplexing (SEFDM), which enhances spectral utilization by employing non-orthogonal overlapped subcarriers in the frequency domain [20, 21]. This enables the transmission of more symbols within a given bandwidth compared to conventional OFDM systems. To support practical implementation, an efficient inverse DFT (IDFT)-based method has been proposed in [22] to generate SEFDM signals with complexity similar to that of OFDM. However, the orthogonality loss in SEFDM introduces inter-carrier interference (ICI), which has been characterized through mathematical analysis and simulation modeling in [23]. Linear detectors such as zero forcing [24] and MMSE [25] have been investigated for their simplicity, but suffer from degraded error rate performance under strong ICI. To improve performance of MMSE detector, an iterative detection (ID) method has been proposed in [26] to progressively mitigate interference. Nonetheless, as a non-orthogonal extension of OFDM, SEFDM inherits sensitivity to Doppler effects, leading to notable performance degradation under rapid channel variations and limiting its applicability in real-time high-mobility scenarios.
@@ I-E Notation
Bold lowercase and uppercase letters denote vectors and matrices, respectively. $\mathbb{C}^{M\times N}$ represents the space of complex-valued matrices of size $M\times N$. $\operatorname{diag}(\mathbf{a})$ denotes a diagonal matrix with the elements of vector $\mathbf{a}$ on its main diagonal. The operators $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ denote the complex conjugate, transpose, Hermitian, and matrix inverse, respectively. ${\bf{I}}_{M}$ denotes the $M\times M$ identity matrix. $\delta(\cdot)$ represents the Dirac delta function, and $(\cdot)_{N}$ denotes the modulo-$N$ operation. $\Re{\cdot}$ and $\Im{\cdot}$ denote the real and imaginary parts of a complex quantity, respectively. The expectation operator is denoted by $\mathbb{E}\{\cdot\}$.
@@ II-A Modulation
where $\mathbf{A}_{\alpha}=\mathbf{\Lambda}_{c_{2}}\mathbf{F}_{\alpha}\mathbf{\Lambda}_{c_{1}}$ represents the modulation matrix for nAFDM, with $\mathbf{\Lambda}_{c}=\text{diag}\left(e^{-\imath 2\pi cn^{2}},\,n=0,\ldots,N-1\right)$ and $\mathbf{F}_{\alpha}$ having entries of $\frac{1}{\sqrt{N}}e^{-\imath 2\pi\frac{\alpha mn}{N}}$.
@@
$s_{\alpha}[n+kN]$ $=\frac{1}{\sqrt{N}}e^{\imath 2\pi c_{1}(k^{2}N^{2}+2kNn)}$
$\times\sum_{m=0}^{N-1}e^{\imath 2\pi\alpha mk}\cdot x[m]e^{\imath 2\pi\left(c_{1}n^{2}+c_{2}m^{2}+\frac{\alpha nm}{N}\right)},$ (3)
where $k\in\mathbb{Z}$ denotes the period index. Due to the presence of the term $e^{\imath 2\pi\alpha mk}\neq 1$ when $\alpha\in(0,1)$, (II-B) cannot be simplified to a phase-shifted version of $s_{\alpha}[n]$, i.e., $s_{\alpha}[n+kN]\neq\frac{1}{\sqrt{N}}e^{\imath 2\pi c_{1}(k^{2}N^{2}+2kNn)}s_{\alpha}[n]$, which indicates that the chirp-periodic structure is no longer preserved, thereby rendering the CPP approach is incompatible. To address this issue, the proposed nAFDM adopts a conventional CP instead of a CPP, which is defined as
@@ II-C Channel Model
The channel impulse response of the $P$-path doubly selective channel is expressed as
@@
where $h_{i}$, $l_{i}$, and $f_{i}$ denote the channel fading coefficients, integer delay shifts, and Doppler shifts (in digital frequencies) of the $i$-th path, respectively. Let $\nu_{i}\triangleq Nf_{i}=a_{i}+\beta_{i}$ represent the Doppler shift normalized with respect to the subcarrier spacing, where $a_{i}\in[-a_{\max},a_{\max}]$ and $\beta_{i}\in\left(-\frac{1}{2},\frac{1}{2}\right]$ denote the integer and fractional parts of $\nu_{i}$, respectively. Note that $\nu_{\max}\triangleq{\text{max}}(\nu_{i})$ and $l_{\max}\triangleq{\text{max}}(l_{i})$ represent the normalized maximum Doppler shift and the maximum delay, respectively.
@@ II-E Input-Output Relationship
Fig. 2 shows the structure of the channel matrix $\mathbf{H}_{i}$ under different Doppler conditions and compression factors. For conventional AFDM, when Doppler shifts are integers, each row or column of $\mathbf{H}_{i}$ contains exactly one non-zero entry. On the other hand, for non-integer Doppler shifts, a small cluster of secondary non-zero entries appears adjacent to the main path due to fractional Doppler leakage. In contrast, in nAFDM, due to the non-integer nature of $\alpha$, it becomes difficult for $p$ and $q$ to precisely satisfy the integer alignment condition, causing the main path energy to split and spread across multiple adjacent positions. Furthermore, the distribution pattern of $\mathbf{H}_{i}$ under fractional Doppler shifts is similar to the integer Doppler shift case, as the combined effect of the Doppler shift and bandwidth compression factor introduces fractional phase components even for integer Doppler values. As a result, the energy in the effective channel matrix becomes more dispersed, with multiple non-zero elements emerging beyond the main peak and spreading across almost the entire matrix. Additionally, we observe that some rows or columns of the $\mathbf{H}_{i}$ only contain very small values. This is due to the non-orthogonality introduced by bandwidth compression, which causes the ideal peak location $q^{*}$ to be non-integer, or even fall outside the allowable index range. In other words, for certain rows, no integer column index $q\in[0,N-1]$ enables $H_{i}[p,q]$ to reach its ideal peak value.
@@ III-A Signal Generation
Specifically, we employ an IDFT of length $N^{\prime}=\frac{N}{\alpha}$. The new input symbol vector ${\bf{x^{\prime}}}\in\mathbb{C}^{N^{\prime}\times 1}$ is defined as
@@
Notably, the proposed nAFDM framework can be configured to implement conventional AFDM, OFDM, OCDM, and SEFDM systems by appropriately selecting the parameters $c_{1}$, $c_{2}$, and $\alpha$. Specifically, setting $\alpha=1$ and $c_{1}=\frac{2(\nu_{\max}+\xi_{\nu})+1}{2N}$ yields the orthogonal AFDM system, where the spacing factor $\xi_{\nu}$ is used to mitigate fractional Doppler effects [5]; setting $\alpha=1$ and $c_{1}=c_{2}=0$ yields the orthogonal OFDM system; setting $\alpha=1$ and $c_{1}=c_{2}=\frac{1}{2N}$ yields the orthogonal OCDM system; and setting $\alpha\in(0,1)$ and $c_{1}=c_{2}=0$ yields the non-orthogonal SEFDM system.
@@ IV-C Soft ID-Assisted ICI Cancellation
From (26), it can be seen that the signal $\bf{\bar{x}}$ is the distorted transmitted signal $\bf{x}$ due to the ICI matrix $\bf{C}_{\alpha}$. Inspired by [26, 28], we propose a soft ID method to mitigate ICI and reconstruct the original transmitted signal, as illustrated in Fig. 6. At each iteration, the algorithm computes symbol probabilities and updates the estimates, progressively mitigating ICI. The main steps of the proposed soft ID method are detailed as follows:

### v2

@@ Abstract
This paper proposes a novel non-orthogonal affine frequency division multiplexing (nAFDM) waveform for reliable high-mobility communications with enhanced spectral efficiency (SE). The key idea is to introduce a bandwidth compression factor into the AFDM modulator to enable controllable subcarrier overlapping. We first detail the proposed nAFDM transceiver and derive the corresponding input-output signal relationship. Then, an efficient nAFDM signal generation method based on the inverse discrete Fourier transform (IDFT) is proposed, enabling practical implementation using existing inverse fast Fourier transform (IFFT) modules without additional hardware complexity. Next, to characterize the impact of non-orthogonal modulation, we derive a closed-form expression of inter-carrier interference (ICI), showing its dependence on the bandwidth compression factor. To mitigate the resulting interference, we propose a soft iterative detection algorithm and a low-complexity implementation approach that leverages the distribution characteristics of ICI. Simulation results demonstrate that 1) in terms of bit error rate (BER), the proposed nAFDM can achieve near identical BER compared to conventional AFDM, while outperforms other waveform counterparts; 2) nAFDM is capable of striking higher SE compared to other existing waveforms; and 3) the proposed nAFDM achieves an attractive BER vs. SE trade-off, and the proposed soft iterative detection (ID) scheme can attain a trade-off between BER and complexity.
@@ I-A Background
Providing reliable wireless connectivity in high-mobility scenarios, such as vehicle-to-everything systems, high-speed railways, unmanned aerial vehicles, and low-earth-orbit satellite networks, is crucial for the upcoming sixth generation systems [1, 2, 3]. The widely adopted orthogonal frequency division multiplexing (OFDM) systems may suffer from severe inter-carrier interference (ICI) in high-mobility environments due to large Doppler shifts, leading to significant degradation in communication performance [4, 5, 6]. Against this problem, a two-dimensional modulation waveform called orthogonal time frequency space (OTFS) has been proposed to improve reliability in high-mobility channels, where information symbols are first multiplexed onto the delay-Doppler grids and then transformed to the time-frequency domain via the inverse symplectic finite Fourier transform (ISFFT) [7]. In addition, OTFS has been extensively investigated and further extended by integrating advanced transmission paradigms to improve spectral efficiency (SE), such as multiple-input multiple-output (MIMO) designs [8], spatial modulation-aided schemes [9], and OTFS-based multiple access frameworks [10]. Unlike OTFS which involves two-dimensional delay-Doppler modulation, affine frequency division multiplexing (AFDM) emerges recently as a promising backwards-compatible multicarrier waveform in [11]. In AFDM, data symbols are modulated over multiple orthogonal chirp subcarriers, thus spreading each information symbol across the entire bandwidth [12]. Such a modulation is carried out via inverse discrete affine Fourier transform (IDAFT) [13], which can be efficiently implemented with the aid of inverse fast Fourier transform (IFFT) [11]. At the receiver side, DAFT is applied to recover the transmitted symbols. Since DAFT can be regarded as a generalization of both the discrete Fourier transform (DFT) and the discrete Fresnel transform (DFnT) [14], AFDM can be built upon the minimum modification of OFDM systems and includes orthogonal chirp division multiplexing (OCDM) as a special case [15]. Moreover, AFDM can be cast as a DFT-based precoded-OFDM waveform [16] and was incorporated into a unified single-carrier interleaved frequency-division multiplexing (SC-IFDM)-based transceiver, enabling orthogonal coexistence and backwards-compatible implementations within existing FFT/IFFT modules [17]. Compared to OFDM and OCDM, AFDM enables separable channel representation and achieves full diversity in doubly selective channels by appropriately tuning the chirp rate to match the Doppler profile [11, 18]. These advantages make AFDM an attractive waveform for achieving enhanced communication performance in high-mobility environments.
@@ I-B Related Works
Several detection and channel estimation algorithms have been proposed to enhance the performance of AFDM. These include a maximum ratio combining scheme [11, 19], an iterative minimum mean square error (MMSE) receiver that balances complexity and detection reliability [18], and a message passing algorithm that exploits the sparse structure of the DAFT-domain channel to jointly perform interference cancellation and signal detection with reduced complexity [20]. Pilot-aided channel estimation for AFDM systems was first introduced in [11], where pilot symbols are embedded into the transmitted frame to efficiently estimate channel parameters. This approach was later extended to MIMO-AFDM systems in [21], where a low-complexity diagonal reconstruction algorithm was developed to estimate the effective channel matrix. To enhance SE, a channel estimation scheme leveraging superimposed pilots in the DAFT domain was proposed [22].
Besides, AFDM has been integrated with other advanced transmission techniques for a wide range of applications. For example, it was introduced to serve as the foundational multicarrier waveform to support massive high-mobility machine-type communications through the integration of sparse code multiple access [23]. In addition, AFDM with index modulation [24, 25] and generalized spatial modulation aided AFDM [26] were proposed for improved bit error rate (BER) performance and energy efficiency. Furthermore, AFDM has also been investigated in the context of integrated sensing and communications (ISAC). In [27], an AFDM-based ISAC system was developed and it was shown that accurate range-velocity sensing can be achieved using only a single DAFT-domain pilot. The ambiguity function properties of AFDM signals for range-velocity sensing were later analyzed in [28], and two performance metrics for AFDM-ISAC, namely sensing SE and sensing outage probability, were introduced in [29]. In [30], a Bayesian parametric bilinear Gaussian belief propagation framework was proposed for AFDM-ISAC to enable joint channel estimation, data detection, and radar-parameter estimation. In addition, recent studies have explored physical-layer security for AFDM transmissions by exploiting the flexibility of chirp parameters, including secret pre-chirp sequence permutation [31] and reciprocity-based user-specific pre-chirp parameter generation [32]. Several modified affine-waveform variants have been proposed to improve AFDM characteristics. In particular, affine filter bank modulation was proposed in [33, 34], which integrates filter bank multicarrier processing into the AFDM framework to reduce the peak-to-average power ratio and out-of-band emission. Chirp-permuted AFDM was introduced in [35] by incorporating a chirp-permutation domain on top of conventional AFDM, thereby enhancing ambiguity function resolution and the peak-to-sidelobe ratio in the Doppler domain.
@@
To meet the ever-growing demand for higher SE, non-orthogonal multicarrier waveform designs have attracted increasing research attention. One notable example is spectrally efficient frequency division multiplexing (SEFDM), which enhances spectral utilization by employing non-orthogonal overlapped subcarriers in the frequency domain [36, 37]. This enables the transmission of more symbols within a given bandwidth compared to conventional OFDM systems. To support practical implementation, an efficient inverse DFT (IDFT)-based method was proposed in [38] to generate SEFDM signals with complexity similar to that of OFDM. However, the orthogonality loss in SEFDM introduces inter-carrier interference (ICI), which was characterized through mathematical analysis and simulation modeling in [39]. Linear detectors such as zero forcing [40] and MMSE [41] were investigated for their simplicity, but suffer from degraded error rate performance under strong ICI. To improve the performance of the MMSE detector, an iterative detection (ID) method was proposed in [42] to progressively mitigate interference. Nonetheless, as a non-orthogonal extension of OFDM, SEFDM inherits sensitivity to Doppler effects, leading to notable performance degradation under rapid channel variations and limiting its applicability in real-time high-mobility scenarios.
@@ I-E Notation
Bold lowercase and uppercase letters denote vectors and matrices, respectively. $\mathbb{C}^{M\times N}$ represents the space of complex-valued matrices of size $M\times N$. $\operatorname{diag}(\mathbf{a})$ denotes a diagonal matrix with the elements of vector $\mathbf{a}$ on its main diagonal. The operators $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{-1}$ denote the complex conjugate, transpose, Hermitian, and matrix inverse, respectively. ${\bf{I}}_{M}$ denotes the $M\times M$ identity matrix. $\delta(\cdot)$ represents the Dirac delta function, and $(\cdot)_{N}$ denotes the modulo-$N$ operation. $\Re\{\cdot\}$ and $\Im\{\cdot\}$ denote the real and imaginary parts of a complex quantity, respectively. The expectation operator is denoted by $\mathbb{E}\{\cdot\}$.
@@ II-A Modulation
where $\mathbf{A}_{\alpha}=\mathbf{\Lambda}_{c_{2}}\mathbf{F}_{\alpha}\mathbf{\Lambda}_{c_{1}}$ represents the modulation matrix for nAFDM, with $\mathbf{\Lambda}_{c}=\text{diag}\left(e^{-\imath 2\pi cn^{2}},\,n=0,\ldots,N-1\right)$ and $\mathbf{F}_{\alpha}$ having entries of $\frac{1}{\sqrt{N}}e^{-\imath 2\pi\frac{\alpha mn}{N}}$. Moreover, $\mathbf{F}_{\alpha}$ can be viewed as the scaled principal $N\times N$ submatrix of an $N^{\prime}$-point DFT matrix $\bar{\mathbf{F}}\in\mathbb{C}^{N^{\prime}\times N^{\prime}}$ with $N^{\prime}=N/\alpha$, see (58) in Appendix B.
@@
$s_{\alpha}[n+\tau N]$ $=\frac{1}{\sqrt{N}}e^{\imath 2\pi c_{1}(\tau^{2}N^{2}+2\tau Nn)}$
$\times\sum_{m=0}^{N-1}e^{\imath 2\pi\alpha m\tau}\cdot x[m]e^{\imath 2\pi\left(c_{1}n^{2}+c_{2}m^{2}+\frac{\alpha nm}{N}\right)},$ (3)
where $\tau\in\mathbb{Z}$ denotes the period index. Due to the presence of the term $e^{\imath 2\pi\alpha m\tau}\neq 1$ when $\alpha\in(0,1)$, (II-B) cannot be simplified to a phase-shifted version of $s_{\alpha}[n]$, i.e., $s_{\alpha}[n+\tau N]\neq e^{\imath 2\pi c_{1}(\tau^{2}N^{2}+2\tau Nn)}s_{\alpha}[n]$, which indicates that the chirp-periodic structure is no longer preserved, thereby rendering the CPP approach is incompatible. To address this issue, the proposed nAFDM adopts a conventional CP instead of a CPP, which is defined as
@@ II-C Channel Model
The channel impulse response of the $P$-path doubly selective channel at time $n$ and delay $l$ is expressed as
@@
where $h_{i}$, $l_{i}$, and $f_{i}$ denote the channel fading coefficients, integer delay shifts, and Doppler shifts (in digital frequencies) of the $i$-th path, respectively. Let $\nu_{i}\triangleq Nf_{i}=a_{i}+\beta_{i}$ represent the Doppler shift normalized with respect to the subcarrier spacing, where $a_{i}\in\mathbb{Z}$ is the integer part of $\nu_{i}$ with $a_{i}\in[-a_{\max},a_{\max}]$, and $\beta_{i}\in\left(-\frac{1}{2},\frac{1}{2}\right]$ denotes the fractional part. Note that $\nu_{\max}\triangleq{\text{max}}(\nu_{i})$ and $l_{\max}\triangleq{\text{max}}(l_{i})$ represent the normalized maximum Doppler shift and the maximum delay, respectively.
@@ II-E Input-Output Relationship
Fig. 2 shows the structure of the channel matrix $\mathbf{H}_{i}$ under different Doppler conditions and compression factors. For conventional AFDM, when Doppler shifts are integers, each row or column of $\mathbf{H}_{i}$ contains exactly one non-zero entry. On the other hand, for non-integer Doppler shifts, a small cluster of secondary non-zero entries appears adjacent to the main path due to fractional Doppler leakage. In contrast, in nAFDM, due to the non-integer nature of $\alpha$, it becomes difficult for $p$ and $q$ to precisely satisfy the integer alignment condition, causing the main path energy to split and spread across multiple adjacent positions. Furthermore, the distribution pattern of $\mathbf{H}_{i}$ under integer Doppler shifts exhibits a spreading behavior similar to that in the fractional Doppler case, as the combined effect of the Doppler shift and the bandwidth compression factor introduces fractional phase components even for integer Doppler values. As a result, the energy in the effective channel matrix becomes more dispersed, with multiple non-zero elements emerging beyond the main peak and spreading across almost the entire matrix. Additionally, we observe that some rows or columns of the $\mathbf{H}_{i}$ only contain very small values. This is due to the non-orthogonality introduced by bandwidth compression, which causes the ideal peak location $q^{*}$ to be non-integer, or even fall outside the allowable index range. In other words, for certain rows, no integer column index $q\in[0,N-1]$ enables $H_{i}[p,q]$ to reach its ideal peak value.
@@ III-A Signal Generation
Specifically, we employ an IDFT of length $N^{\prime}=\frac{N}{\alpha}$, where $\alpha\in(0,1)$ is selected such that $N^{\prime}\in\mathbb{N}$ for practical FFT/IFFT implementation. The new input symbol vector ${\bf{x^{\prime}}}\in\mathbb{C}^{N^{\prime}\times 1}$ is defined as
@@
Proof: The detailed derivation of $s_{\alpha}$ is given in Appendix B. $\blacksquare$
Notably, the proposed nAFDM framework can be configured to implement conventional AFDM, OFDM, OCDM, and SEFDM systems by appropriately selecting the parameters $c_{1}$, $c_{2}$, and $\alpha$. Specifically, setting $\alpha=1$ and $c_{1}=\frac{2(\alpha_{\max}+\xi_{\nu})+1}{2N}$ yields the orthogonal AFDM system, where the spacing factor $\xi_{\nu}$ is used to mitigate fractional Doppler effects [11]; setting $\alpha=1$ and $c_{1}=c_{2}=0$ yields the orthogonal OFDM system; setting $\alpha=1$ and $c_{1}=c_{2}=\frac{1}{2N}$ yields the orthogonal OCDM system; and setting $\alpha\in(0,1)$ and $c_{1}=c_{2}=0$ yields the non-orthogonal SEFDM system.
@@ IV-C Soft ID-Assisted ICI Cancellation
From (26), it can be seen that the signal $\bf{\bar{x}}$ is the distorted transmitted signal $\bf{x}$ due to the ICI matrix $\bf{C}_{\alpha}$. Inspired by [42, 44], we propose a soft ID method to mitigate ICI and reconstruct the original transmitted signal, as illustrated in Fig. 6. Specifically, the proposed method exploits symbol posterior probabilities to update the soft symbol estimates, which are then used for progressive interference cancellation. Moreover, a selective redetection step is performed for certain selected low-reliability symbols according to their estimated variances and posterior probabilities, providing an additional refinement to the final decisions. The main steps of the proposed soft ID method are detailed as follows: